# Investigating the Impact of Project Risks on Employee Turnover Intentions in the IT Industry of Pakistan

### v3

@@ 3.5 Organizational Context/ Macro Level
Researchers have introduced additional organizational context variables such as unit-level attitudes, engagement, satisfaction, and climate approach. Employee engagement and satisfaction across various business units are not positively associated with employee turnover Harter and Harter (2002). Furthermore,  McElroy JC1 (2001) have reported that the performance of a unit is negatively affected by both involuntary and voluntary turnover. When controlling for the other two schemes, each scheme has negative consequences, albeit with different effects.  Koys (2001) has stated that citizenship and level of unit satisfaction behavior from the first year positively forecast satisfaction of customers and unit profit in the second year. On the other hand, turnover from the first year outcome has negatively predicted second-year results. Kacmar (2006) has shown that turnover at the unit level is negatively linked with unit performance. This connection was mediated by unit efficiency, indicating that workforce stability is required for organizational effectiveness.
Researchers have also examined the impact of organizational diversity on employee turnover. Elvira and Cohen investigated the effect of gender composition on one’s own sex at various positions within the organization. Their results indicated that employees of the same sex had an impact on the retention of female employees, negatively affecting the turnover rate, but had no impact on male employees. Research on public accountants conducted by  Saroj Parasuraman (1992) reported that female public accountants are more likely to leave their positions compared to their male counterparts, not due to family issues but because of the lack of promotional opportunities. Conversely, research results by Lyness (2001) showed that female financial managers were less likely to leave compared to male financial managers when they obtained a higher position in the preceding eleven months. Demographic misfit within work groups is also associated with a high turnover hazard. A comprehensive study conducted by Mallol (2007) on more than 20 US corporations and 450,000 professionals and managers found that incumbents of jobs typically held by African Americans and Hispanics were at a higher risk of turnover. Other organizational variables, such as firms’ pay distribution, also affect employee turnover Matt and John (2002).  Eisenberger R1 (2002) found that the negative effect of perceived supervisory support on employee turnover is mediated by perceived organizational support. However, a study by  753 (a) showed that perceived supervisor support has an independent effect on turnover cognition, and perceived organizational support had a significant impact on turnover mediated through affective and organizational commitment.
@@ 3.6 Person Context Interface
Current research in this perspective focuses extensively on interpersonal relations and the interface between employees and their environment, a dimension largely overlooked by previous models from the 1990s era. In recent studies, turnover has been examined from a relational perspective. For instance, Friedman and Holtom (2002a) studied the impact of minority network associations on minority employee turnover, highlighting the significance of social embeddedness in understanding turnover dynamics. Similarly, research by  Colquitt JA1 (2001) Hendrix (1998) Tekleab et al. (2005) emphasizes the importance of justice perceptions in understanding satisfaction and commitment among employees. Sequential linkages and the significance of evidence from procedural and interactional justice to employee turnover and commitment have been highlighted by  Simons (2003). They reported that Person-Organization fit predicts turnover, partially mediated by job cognition and attitudes. Additionally,  Wright (2007) found that psychological well-being moderates the association between satisfaction and turnover, with a negative relationship observed for employees with lower levels of personal well-being, while this relationship was less significant for those with higher levels of well-being. Contrary to expectations, both leader influence between socialization and perceived supervisor support and social citizenship behaviors were not positively related to turnover. Another research study suggested that employees with lower levels of supervisor-rated behaviors, despite displaying organizational citizenship, have a higher likelihood of quitting their jobs Chen et al. (1998).
@@ 3.7 Organizational Level Consequences
There are several organizational-level consequences associated with employee turnover. Shaw and colleagues have expanded on the previous perspective of social capital, suggesting that the impact on performance resulting from employee turnover is more complex than simply the loss of human capital  Dess (2001). They have demonstrated that the loss of organizational capital due to turnover is negatively correlated with efficiency, and the level of organizational turnover moderates this relationship. Shaw, Gupta, et al. found a curve linear relationship of workforce performance and turnover in a two-unit turnover study. The negative impact was stronger when turnover was low but weakened when turnover was high.
Mitchell (2001) introduced the concept of job embeddedness, which focuses on the various reasons that influence individuals to stay in their jobs. The main aspects of job embeddedness include the connections employees have with others within the organization and the community, how well they fit within the organization, and what they would have to sacrifice if they were to leave. These three dimensions of job embeddedness are referred to as links, fit, and sacrifice, and they are relevant in both the organizational and community contexts. Mitchell et al. reported the following results after conducting experiments on samples of retail and hospital employees.
@@
Lee, Mitchel, Sablynski, Burton, and Holtom further expanded on research regarding job embeddedness. Their significant contribution involved disaggregating job embeddedness into two major sub-dimensions. Utilizing a large sample of bank employees, their results demonstrated that off-the-job embeddedness could significantly predict voluntary turnover. Additionally, the research explored potential differences between Caucasian and Hispanic groups concerning turnover and job embeddedness Mallol (2007). Their findings indicated that job embeddedness may vary across demographic groups but does not strongly predict employee retention.
@@
The research conducted by  Allen and O’Brien (2006) has revealed that organizational social tactics such as collective, fixed, and investiture methods facilitate the active integration of new employees into the organization. Their findings indicate that job embeddedness acts as a mediator in the relationship between socialization tactics and employee turnover.
@@
Furthermore,  Kevin et al. (2005) discovered that network centrality and interpersonal citizenship behavior were negatively correlated with turnover, suggesting that employees with stronger networks and better interpersonal relationships are less likely to leave their jobs. Additionally,  Zatzick and Iverson (2006) published findings suggesting that off-the-job embeddedness reduces turnover among women, highlighting the importance of factors outside of the workplace in influencing turnover decisions.
@@
Moreover, Crossley (2007) undertook a reconceptualization of job embeddedness, transitioning it from a formative to a reflective construct. They investigated how general and composite forms of job embeddedness converge into Mobley Type turnover parameters. Their results demonstrated that general job embeddedness was strongly associated with intentions to quit, searching for new job opportunities, and actual turnover. Conversely, composite job embeddedness was linked to the intention to search for and quit a job but not to actual turnover.
@@
Holtom and Inderrieden (2006) has endeavored to unify the parameters of job embeddedness theory and develop a comprehensive model. Their experimental findings from a study conducted at the national level, focusing on employees staying and leaving, indicated that job stayers exhibit the highest level of job embeddedness. Following them, shock-induced leavers showed the next highest level of job embeddedness, while non-shock-induced leavers demonstrated the lowest level. Additionally,  David et al. (2003) found that supportive human resource practices in organizations, such as fairness of rewards and participation in decision-making, increase perceived organizational support (POS) among employees.
@@
Furthermore,  Mobley (1982) has advocated for further research in the area of dynamic modeling of the turnover process, emphasizing how turnover unfolds over time. Many researchers have delved into dynamically modeling the turnover process, incorporating behavioral and attitudinal changes over time. For instance, researchers like Sturman and Trevor discovered that quitters’ performance remains relatively stable over time, whereas stayers’ performance shows positive trends.
@@
Steel (2002) has introduced a novel job searching model, delineating three distinct job phases—passive planning, focused searching, and communicating with potential employers—and two gateways of job searching: instant job offers and financial benefits. They underscore the importance of knowledge acquisition during the job search phase, asserting that dynamic search processes enhance individuals’ understanding of the employment market, especially for those contemplating quitting their current job. Moreover, they highlight that the relationship between perceived job alternatives and actual outcomes differs between individuals who intend to stay at their current organization and those who are less advanced in their job search process.
@@
Moreover,  Bentein (2005) demonstrated that changes in effective and normative assurance are associated with variations in turnover intention over a six-month period. Changes in turnover intention over three months are related to subsequent changes over the following three months.
@@ 3.8 Employee Turnover In Context With IT Industry
For example,  Christina Ling-Hsing Changa (2012); March and Simon (1958); 75323:2894206; Wuensch (2005) conducted research on the impact of social networks and value connectivity on turnover intentions among employees in both public and private sectors. They found that strong intra-organizational social networks, characterized by good relationships and a sense of responsibility towards colleagues, reduce turnover intentions. Conversely, active external social networks increase external opportunities, making it more likely for employees to consider leaving the organization. They also proposed that person-organization fit plays a role in shaping turnover intentions, suggesting that employees with a strong fit in terms of value congruence are more likely to exhibit long-term commitment.
@@
Similarly,  Christina Ling-Hsing Changa (2012) identified various factors that disrupt the relationship between job satisfaction and the intention to seek employment elsewhere. Through qualitative analysis of 10 cases, they identified new disruptions, contributing to a deeper understanding of employee turnover patterns.
@@
According to Damien Joseph and Ang. (2007), individuals working in the IT industry operate in a dynamic environment that necessitates regular skill renewal. They argue that an individual’s growth and development needs are crucial factors influencing turnover intentions. Their model suggests that the strength of development needs interacts with job satisfaction to influence turnover intentions. Additionally, factors such as the motivational aspects of the job, role conflict, and role ambiguity affect turnover intentions through their impact on job satisfaction.
@@
Other researchers have explored the transformative effects of leadership on creativity within organizations. For instance, research by  Yıldız (2014) has demonstrated that transformational leaders enhance employees’ performance expectations and creativity levels. They distinguish between individual, organizational, and group creativity levels and highlight the significant role of organizational climate in fostering creativity. They also find that the link between transformational leadership and creativity is moderated by organizational climate. There statistics suggest that a large number of IT professionals are expected to retire in the next decade, while fewer new workers are entering the IT field. This is expected to increase the demand for skilled personnel in the industry, making it more challenging for organizations to retain their employees. Considering this,  Jordan Shropshire (2012) have identified reasons that lead IT workers to change fields, proposing contextual frameworks involving variables such as work pressure, job insecurity, and burnout. They also suggest that age moderates each of these paths.
@@
Personal desires and goals play a significant role in the IT industry. When employees’ goals and desires are not met, they become dissatisfied, affecting their motivation and other factors contributing to employee turnover. Career anchors represent the personal aspirations of employees, crucial for organizations to attract and retain top talent. Career stages illustrate employees’ advancement paths within an organization, reflecting their professional growth and value to the organization.
@@
Furthermore, research by Christina Ling-Hsing Chang, Victor Chen, Gary Klein, and James J. Jiang has examined the career transitions of 10 IS employees, noting changes in career anchors throughout their careers. They found that technical proficiency and security anchors remain significant at all career stages, while factors such as geographic security, managerial competence, and self-determination become more significant in later stages.
@@
Moreover,  Arpita Sharmaa (2012) have highlighted that software development in organizations is often constrained by various risks at both micro and macro levels. They identify mitigation strategies proposed by numerous workers and emphasize how the organizational environment influences individuals’ abilities to identify and address software-related risks. Their research identifies dimensions of organizational environment and risk within the software industry in India through factor analysis.
@@ 4 Theoretical Model And Hypothesis
We’ve constructed a theoretical framework that reflects the findings of the reviewed literature for this research endeavor.The developed conceptual framework is in line with the objectives of our study, visually depicted in Figure 1.This research framework, illustrated graphically, outlines the independent variables: project-specific risk factors, internal and external networking, and external social network.Job satisfaction serves as a mediating variable, while perceived job alternatives function as the moderating variable for the dependent variable, turnover intentions.Throughout our investigation, we have explored several hypotheses.
@@ 5.2 Sampling Definition
A sample is a subset that captures the characteristics of a larger population  Peck (2008). When dealing with large populations, samples used for statistical sampling must be unbiased and representative of the larger population. Defining criteria for sample size is crucial in ensuring the validity of research findings. In this study, we have established the following criteria for determining the sample size:
@@ 5.3 Design Of Questionnaire
We have developed a questionnaire based on a comprehensive literature review, selecting recent scales to ensure relevance. The questions are designed to be clear and direct, facilitating easy understanding for respondents. The questionnaire comprises the following sections:
@@
  • Perceived Job Alternatives: These questions explore how perceived job alternatives mediate the relationship between organizational satisfaction and turnover intention.
@@ 5.4 Survey Procedure
Various data collection techniques exist, such as face-to-face, telephone interviews, and questionnaires.In our research, we opted to distribute a questionnaire via the web.While this method may exclude individuals without computer access, we focused on IT professionals who are likely to have computer access.Therefore, we distributed our questionnaire online throughout the IT industry in major cities across Pakistan.
@@ 6 Data Analysis
We conducted various analyses to validate our hypotheses. At first, we examined Pearson’s correlation to assess the relationships between:
@@
In the second section, we performed regression analyses to determine the impact of:

### v4

@@ 3.5 Organizational Context/ Macro Level
Researchers have introduced additional organizational context variables such as unit-level attitudes, engagement, satisfaction, and climate approach. Employee engagement and satisfaction across various business units are not positively associated with employee turnover 75323:2894227. Furthermore,  75323:2894258 have reported that the performance of a unit is negatively affected by both involuntary and voluntary turnover. When controlling for the other two schemes, each scheme has negative consequences, albeit with different effects.  75323:2894228 has stated that citizenship and level of unit satisfaction behavior from the first year positively forecast satisfaction of customers and unit profit in the second year. On the other hand, turnover from the first year outcome has negatively predicted second-year results. 75323:2894211 has shown that turnover at the unit level is negatively linked with unit performance. This connection was mediated by unit efficiency, indicating that workforce stability is required for organizational outcomes.
Researchers have also examined the impact of organizational diversity on employee turnover. Elvira and Cohen investigated the effect of gender composition on one’s own gender at various positions within the organization. Their results indicated that employees of the same sex had an impact on the retention of female employees, negatively affecting the turnover rate, but had no impact on male employees. Research on public accountants conducted by  75323:2894274 reported that female public accountants are more likely to leave their positions compared to their male counterparts, not due to family issues but because of the lack of promotional opportunities. Conversely, research results by 75323:2894229 showed that female financial managers were less likely to leave compared to male financial managers when they obtained a higher position in the preceding eleven months. Demographic misfit within work groups is also associated with a high turnover hazard. A comprehensive study conducted by 75323:2894237 on more than 20 US corporations and 450,000 professionals and managers found that incumbents of jobs typically held by African Americans and Hispanics were at a higher risk of turnover. Other organizational variables, such as firms’ pay distribution, also affect employee turnover 75323:2894253.  75323:2894259 found that the negative effect of perceived supervisory support on employee turnover is mediated by perceived organizational support. However, a study by  75323:2894231 showed that perceived supervisor support has an independent effect on turnover cognition, and perceived organizational support had a notable impact on turnover mediated through affective and organizational commitment.
@@ 3.6 Person Context Interface
Current research in this perspective focuses extensively on interpersonal relations and the interface between employees and their environment, a dimension largely overlooked by previous models from the 1990s era. In recent studies, turnover has been examined from a relational perspective. For instance, 75323:2894251 studied the impact of minority network associations on minority employee turnover, highlighting the significance of social embeddedness in understanding turnover dynamics. Similarly, research by  75323:2894260 75323:2894233 75323:2894261 emphasizes the importance of justice perceptions in understanding satisfaction and commitment among employees. Sequential linkages and the significance of evidence from procedural and interactional justice to employee turnover and commitment have been highlighted by  75323:2894234. They reported that Person-Organization fit predicts turnover, partially mediated by job cognition and attitudes. Additionally,  75323:2894256 found that psychological well-being moderates the association between satisfaction and turnover, with a negative relationship observed for employees with lower levels of personal well-being, while this relationship was less notable for those with higher levels of well-being. Contrary to expectations, both leader influence between socialization and perceived supervisor support and social citizenship behaviors were not positively related to turnover. Another research study suggested that employees with lower levels of supervisor-rated behaviors, despite displaying organizational citizenship, have a higher likelihood of quitting their jobs 75323:2894255.
@@ 3.7 Organizational Level Consequences
There are several organizational-level consequences associated with employee turnover. Shaw and colleagues have expanded on the previous perspective of social capital, suggesting that the impact on results resulting from employee turnover is more complex than simply the loss of human capital  75323:2894271. They have showed that the loss of organizational capital due to turnover is negatively correlated with efficiency, and the level of organizational turnover moderates this relationship. Shaw, Gupta, et al. found a curve linear relationship of workforce results and turnover in a two-unit turnover study. The negative impact was stronger when turnover was low but weakened when turnover was high.
75323:2894236 introduced the concept of job embeddedness, which focuses on the various reasons that influence individuals to stay in their jobs. The main aspects of job embeddedness include the connections employees have with others within the organization and the community, how well they fit within the organization, and what they would have to sacrifice if they were to leave. These three dimensions of job embeddedness are referred to as links, fit, and sacrifice, and they are relevant in both the organizational and community contexts. Mitchell et al. reported the following results after conducting experiments on samples of retail and hospital employees.
@@
Lee, Mitchel, Sablynski, Burton, and Holtom further expanded on research regarding job embeddedness. Their useful contribution involved disaggregating job embeddedness into two notable sub-dimensions. Utilizing a large sample of bank employees, their results showed that off-the-job embeddedness could significantly predict voluntary turnover. Additionally, the research explored potential differences between Caucasian and Hispanic groups concerning turnover and job embeddedness 75323:2894237. Their findings indicated that job embeddedness may vary across demographic groups but does not clearly predict employee retention.
@@
The research conducted by  75323:2894273 has revealed that organizational social tactics such as collective, fixed, and investiture methods facilitate the active integration of new employees into the organization. Their findings indicate that job embeddedness acts as a mediator in the relationship between socialization tactics and employee turnover.
@@
Furthermore,  75323:2894270 discovered that network centrality and interpersonal citizenship behavior were negatively correlated with turnover, suggesting that employees with stronger networks and better interpersonal relationships are less likely to leave their jobs. Additionally,  75323:2894263 published findings suggesting that off-the-job embeddedness reduces turnover among women, highlighting the importance of factors outside of the workplace in influencing turnover decisions.
@@
Moreover, 75323:2894272 undertook a reconceptualization of job embeddedness, transitioning it from a formative to a reflective construct. They investigated how general and composite forms of job embeddedness converge into Mobley Type turnover parameters. Their results showed that general job embeddedness was clearly associated with intentions to quit, searching for new job opportunities, and actual turnover. Conversely, composite job embeddedness was linked to the intention to search for and quit a job but not to actual turnover.
@@
75323:2894269 has endeavored to unify the parameters of job embeddedness theory and develop a comprehensive model. Their experimental findings from a study conducted at the national level, focusing on employees staying and leaving, indicated that job stayers exhibit the highest level of job embeddedness. Following them, shock-induced leavers showed the next highest level of job embeddedness, while non-shock-induced leavers showed the lowest level. Additionally,  75323:2894262 found that supportive human resource practices in organizations, such as fairness of rewards and participation in decision-making, increase perceived organizational support (POS) among employees.
@@
Furthermore,  75323:2894204 has advocated for further research in the area of dynamic modeling of the turnover process, emphasizing how turnover unfolds over time. Many researchers have delved into dynamically modeling the turnover process, incorporating behavioral and attitudinal changes over time. For instance, researchers like Sturman and Trevor discovered that quitters’ results remains relatively stable over time, whereas stayers’ results shows positive trends.
@@
75323:2894278 has introduced a relatively unexplored job searching model, delineating three distinct job phases—passive planning, focused searching, and communicating with potential employers—and two gateways of job searching: instant job offers and financial benefits. They underscore the importance of knowledge acquisition during the job search phase, asserting that dynamic search processes enhance individuals’ understanding of the employment market, especially for those contemplating quitting their current job. Moreover, they highlight that the relationship between perceived job alternatives and actual outcomes differs between individuals who intend to stay at their current organization and those who are less advanced in their job search process.
@@
Moreover,  75323:2894264 showed that changes in effective and normative assurance are associated with variations in turnover intention over a six-month period. Changes in turnover intention over three months are related to subsequent changes over the following three months.
@@ 3.8 Employee Turnover In Context With IT Industry
For example,  75323:2894240; 75323:2894201; 75323:2931936 conducted research on the impact of social networks and value connectivity on turnover intentions among employees in both public and private sectors. They found that clear intra-organizational social networks, characterized by good relationships and a sense of responsibility towards colleagues, reduce turnover intentions. Conversely, active external social networks increase external opportunities, making it more likely for employees to consider leaving the organization. They also proposed that person-organization fit plays a role in shaping turnover intentions, suggesting that employees with a clear fit in terms of value congruence are more likely to exhibit long-term commitment.
@@
Similarly,  75323:2894240 identified various factors that disrupt the relationship between job satisfaction and the intention to seek employment elsewhere. Through qualitative analysis of 10 cases, they identified new disruptions, contributing to a deeper understanding of employee turnover patterns.
@@
According to 75323:2894241, individuals working in the IT industry operate in a dynamic environment that necessitates regular skill renewal. They argue that an individual’s growth and development needs are relevant factors influencing turnover intentions. Their model suggests that the strength of development needs interacts with job satisfaction to influence turnover intentions. Additionally, factors such as the motivational aspects of the job, role conflict, and role ambiguity affect turnover intentions through their impact on job satisfaction.
@@
Other researchers have explored the transformative effects of leadership on creativity within organizations. For instance, research by  75323:2894244 has showed that transformational leaders enhance employees’ results expectations and creativity levels. They distinguish between individual, organizational, and group creativity levels and highlight the notable role of organizational climate in fostering creativity. They also find that the link between transformational leadership and creativity is moderated by organizational climate. There statistics suggest that a large number of IT professionals are expected to retire in the next decade, while fewer new workers are entering the IT field. This is expected to increase the demand for skilled personnel in the industry, making it more challenging for organizations to retain their employees. Considering this,  75323:2894245 have identified reasons that lead IT workers to change fields, proposing contextual frameworks involving variables such as work pressure, job insecurity, and burnout. They also suggest that age moderates each of these paths.
@@
Personal desires and goals play a notable role in the IT industry. When employees’ goals and desires are not met, they become dissatisfied, affecting their motivation and other factors contributing to employee turnover. Career anchors represent the personal aspirations of employees, relevant for organizations to attract and retain top talent. Career stages illustrate employees’ advancement paths within an organization, reflecting their professional growth and value to the organization.
@@
Furthermore, research by Christina Ling-Hsing Chang, Victor Chen, Gary Klein, and James J. Jiang has examined the career transitions of 10 IS employees, noting changes in career anchors throughout their careers. They found that technical proficiency and security anchors remain notable at all career stages, while factors such as geographic security, managerial competence, and self-determination become more notable in later stages.
@@
Moreover,  75323:2894243 have highlighted that software development in organizations is often constrained by various risks at both micro and macro levels. They identify mitigation strategies proposed by numerous workers and emphasize how the organizational environment influences individuals’ abilities to identify and address software-related risks. Their research identifies dimensions of organizational environment and risk within the software industry in India through factor analysis.
@@ 4 Theoretical Model And Hypothesis
We’ve constructed a theoretical framework that reflects the findings of the reviewed literature for this research endeavor.The proposed conceptual framework is in line with the objectives of our study, visually depicted in Figure 1.This research framework, illustrated graphically, outlines the independent variables: project-specific risk factors, internal and external networking, and external social network.Job satisfaction serves as a mediating variable, while perceived job alternatives function as the moderating variable for the dependent variable, turnover intentions.Throughout our investigation, we have explored several hypotheses.
@@ 5.2 Sampling Definition
A sample is a subset that captures the characteristics of a larger population  75323:2931894. When dealing with large populations, samples used for statistical sampling must be unbiased and representative of the larger population. Defining criteria for sample size is relevant in ensuring the validity of research findings. In this study, we have explored the following criteria for determining the sample size:
@@ 5.3 Design Of Questionnaire
We have proposed a questionnaire based on a comprehensive literature review, selecting recent scales to aim to relevance. The questions are designed to be clear and direct, facilitating easy understanding for respondents. The questionnaire comprises the following sections:
@@
  • Perceived Job Alternatives: These questions explore how perceived job alternatives moderate the relationship between job satisfaction and turnover intention.
@@ 5.4 Survey Procedure
Various data collection techniques exist, such as face-to-face, telephone interviews, and questionnaires. In our research, we opted to distribute a questionnaire via the web. While this method may exclude individuals without computer access, we focused on IT professionals who are likely to have computer access. Therefore, we distributed our questionnaire online throughout the IT industry in notable cities across Pakistan.
@@ 6 Data Analysis
We conducted various analyses to assess our hypotheses. At first, we examined Pearson’s correlation to assess the relationships between:
@@
In the second section, we performed regression analyses to examine the impact of: